OpenStudy (unklerhaukus):

Solve 2nd Order DE\[y\frac{\text d^2y}{\text dx^2}+1=\left(\frac{\text dy}{\text dx}\right)^2\]

OpenStudy (anonymous):

Interesting method ... $$ y \frac{dv}{dy}v + 1 = v^2 $$

OpenStudy (anonymous):

let dy/dx = v

OpenStudy (anonymous):

let dy/dx=p then d^2y/dx^2=dp/dx=(dp/dy)(dy/dx)=p(dp/dy)

OpenStudy (anonymous):

then we have

OpenStudy (anonymous):

\[yp(dp/dy)+1=p^2\]

OpenStudy (unklerhaukus):

\[y\frac{\text d^2y}{\text dx^2}+1=\left(\frac{\text dy}{\text dx}\right)^2\]\[\text {let } p=\frac{\text dy}{\text dx}\qquad\qquad\frac{\text dp}{\text dx}=\frac{\text d^2y}{\text dx^2}=p\frac{\text dp}{\text dy}\]\[yp \frac{\text dp}{\text dy}+1=p^2\]

OpenStudy (experimentx):

\[ \frac{dp}{dx} = \frac{dp}{dy} \frac{dy}{dx} = \frac{dp}{dy} p\]

OpenStudy (unklerhaukus):

\[yp \frac{\text dp}{\text dy}+1=p^2\]\[yp \frac{\text dp}{\text dy}=p^2-1\]\[\frac{p\text dp}{p^2-1}=\frac{\text dy}{y}\]\[\frac 12 \int\frac{2p}{p^2-1}\cdot \text dp=\int\frac{\text dy}{y}\]

OpenStudy (anonymous):

i made a terrible mistake up there...lol

OpenStudy (unklerhaukus):

\[\frac 12\ln|p^2-1|=\ln|cy|\]\[p^2-1=Ay^2\]\[\left(\frac{\text dy}{\text dx}\right)^2-1=Ay^2\]

OpenStudy (unklerhaukus):

a) have i done this right sofar, b) what can i do next

OpenStudy (anonymous):

I'm not sure if you can do this but, what if you take roots at both sides and; (y')^2=Ay^2 + 1; y'=sqrt(Ay^2+1); separate variables and you have; dy/sqrt(Ay^2+1)=dx; 1/2*argsinh(y)=x + C ???

OpenStudy (foolaroundmath):

@UnkleRhaukus your steps seem correct. Now, \[(\frac{dy}{dx})^{2} = (Ay^{2}+1)\]\[\frac{dy}{dx} = \pm\sqrt{1+Ay^{2}}\] \[\int\limits \frac{dy}{\sqrt{1+Ay^{2}}} = \int\limits \pm dx\]Substitute \[y = \tan\theta\sqrt{A}\] See if you can take it from here.

OpenStudy (turingtest):

I think the sub would be\[y=\frac1{\sqrt A}\tan\theta\]

OpenStudy (turingtest):

...but is that \(\pm\) a wrench in the works? of that I'm not sure

OpenStudy (unklerhaukus):

the answer in the back of the book has logs rather than arctans

OpenStudy (unklerhaukus):

so i think i hav to use this integral for the table \[[20]\qquad\int\frac{\text dt}{a^2-x^2}=\frac{1}{2a}\ln\left|\frac{x+a}{x-a}\right|+c\] somehow

OpenStudy (unklerhaukus):

\[p^2-1=c^2y^2\]